What are the first three terms if the 11th term of an
arithmetic sequence is 60 and the sum of the first 11
terms is 55?

Answer by Edwin McCravy(20056)

What are the first three terms if the 11th term of an
arithmetic sequence is 60 and the sum of the first 11
terms is 55?

a1,__,__,__,__,__,__,__,__,__,60,...

Sn = (n/2)(a1 + an)

S11 = (11/2)(a1 + a11)

55 = (11/2)(a1 + 60)

Multiply both sides by 2

110 = 11(a1 + 60)

110 = 11a1 + 660

-550 = 11a1

-550/11 = a1

-50 = a1 

So

a1,__,__,__,__,__,__,__,__,__,60,

becomes

-50,__,__,__,__,__,__,__,__,__,60,

To fill in the blanks, we only need the common difference d:

an = a1 + (n - 1)d

 60 = -50 + (11 - 1)d
 60 = -50 + 10d
110 = 10d
 11 = d

So the sequence

-50,__,__,__,__,__,__,__,__,__,60,...

becomes

-50,-39,-28,-17,-6,5,16,27,38,49,60,...

So the first three terms are -50,-39 and -28

Edwin